OpenStudy (anonymous):

In parallelogram ABCD, angle A is x(x +3) and angle B is 140°. Find the values of x. my guess is its 7 and 8,let me know if its right

OpenStudy (anonymous):

 angle b is 140 degrees and angle A is x(x+3)

Directrix (directrix):

Angles A and B are supplementary because they are same-side interior angles formed by two parallel lines cut by a transversal. x(x+3) + 140 = 180 x^2 + 3x - 40 = 0 (x - 5) (x + 8) = 0 Zero Product Property x = 5 or x = -8

OpenStudy (anonymous):

iyeah i got that one thank you